OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. a) y = 3x – 6 b) y = 3x – 3 c) y = 3x + 3 d) y = 3x + 6

OpenStudy (anonymous):

m=y2-y1/x2-x1 y=mx+b

OpenStudy (anonymous):

m= -3-6/0-3 Then plug your answer into the slope intercept form

OpenStudy (anonymous):

use this formula \[(y-y _{1})=(y _{2}-y _{1}/x _{2}-x _{1)}(x-x _{1)}\]

OpenStudy (anonymous):